# MergeMix: A Unified Augmentation Paradigm for Visual and Multi-Modal Understanding

### v1

@@ 5.1 Baselines
To evaluate the performance of MergeMix, we compare our MergeMix with some mainstream mixup methods, i.e. Mixup (Zhang et al., 2017), CutMix (Yun et al., 2019), FMix (Harris et al., 2020), SmoothMix (Lee et al., 2020), GridMix (Baek et al., 2021), ResizeMix (Qin et al., 2020), SaliencyMix (Uddin et al., 2021), Attentive-CutMix (Walawalkar et al., 2020), PuzzleMix (Kim et al., 2020), GuidedMixup (Kang & Kim, 2023), AutoMix (Liu et al., 2022) and AdAutoMix (Qin et al., 2024a). DeiT (Touvron et al., 2021), TransMix (Chen et al., 2022), SMMix (Chen et al., 2023) and MixPro (Zhao et al., 2023) for some ViT-based methods. The training configures about datasets and methods follows the open-source library OpenMixup (Li et al., 2022).
@@
To evaluate the training paradigm that we proposed, we compare with three different system-level baselines: (1) SFT with different training paradigms on LLaVA, including LLaVA-NeXT-7/13B (Liu et al., 2024a), SeVa-7B (Zhu et al., 2024), SIMA (Wang et al., 2024), and nSFT (Zhu et al., 2025). (2). Token reduction on LLaVA, including LLaVA-PruMerge+ (Shang et al., 2024), VisionZip (Yang et al., 2025a), VisPrunner (Zhang et al., 2024b), VScan (Zhang et al., 2025a), and LLaVA-Mini (Zhang et al., 2025b). (3). RL training on Qwen2.5-VL-Instruction (Bai et al., 2025), including VisionThink (Yang et al., 2025b).
For all classification results, we report the top-1 test accuracy in the last 10 training epochs for each trial. To facilitate comparison, we mark the best and second best results in bold and cyan. For the LLaVA benchmark, we use the LLaVA official code, and for the Qwen2.5-VL-Instruction benchmark. We use lmms-eval (Zhang et al., 2024a) for evaluation.
Table 1: Top-1 accuracy (%) of mixup methods training 200 epochs on CIFAR100 dataset with different model sizes, T/S/B/L denotes Tiny, Small, Base, and Large, respectively. The full results of training 600 epochs are in Table 9.
@@
Table 2: Top-1 accuracy (%) of mixup methods on the Stanford-Cars dataset. Full results of the CUB200 and FGVC-Aircrafts dataset in Table 10.
@@ 5.2 Datasets
In our paper, we mainly divided into 2 scenarios: Image Classification and MLLM Benchmark. The detailed information about datasets is described in Appendix B.1. For the image classification datasets, we choose 5 public classification datasets, including the small-scale dataset of CIFAR100 (Krizhevsky et al., 2009), the large-scale dataset of ImageNet-1K (Russakovsky et al., 2015), and the fine-grained datasets of CUB200 dataset (Wah et al., 2011), FGVC-Aircrafts dataset (Maji et al., 2013), and Stanford-Cars dataset (Krause et al., 2013). For the MLLM datasets, we choose 16 datasets, including visual question answering (VQAv2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), VizWiz (Gurari et al., 2018), ScienceVQAI (Lu et al., 2022), TextVQA (Singh et al., 2019), MME-RealWorldQA (Zhang et al., 2025c)), understanding (MME (Perception) (Yin et al., 2023), MMBench (Liu et al., 2025), $^{\text{CN}}, MMBench$$^{\text{CC}}$, POPE (F1 score) (Li et al., 2023b), SEEDI (Li et al., 2023a), MMStar (Chen et al., 2024b)), and reasoning (MMMU (Yue et al., 2024a), MMMU-Pro Standard (MMMU-Pros) (Yue et al., 2024b), MathVista (Lu et al., 2024)).
Table 4: Full system-level comparison results in LLaVA. Compared with their counterparts. AVG denotes the average of the nine benchmarks for comprehensive comparison, except for MME, underline denotes MME with the sum of Perception and Cognition. Tokeni denotes training with the token number. Full results in Table 16, Table 17 and Table 18.
@@ 5.3 Implementations
In this subsection, we briefly introduce the implementations on the classification task and the MLLM benchmark. The full description is in Appendix B.2. For the CIFAR100, images are resized to $224\times 224$ for ViT-based models (e.g., DeiT) and trained with AdamW (weight decay 0.05), batch size 100, for 200 or 600 epochs. We use RandomFlip and RandomCrop, plus RandAugment (Cubuk et al., 2020). Learning rates are 1e-3 (DeiT-Tiny/Small, cosine schedule), 5e-4 (ViT-Small/Base), and 2e-4 (ViT-Large). For ImageNet-1K, we adopt the same settings but use a 1e-3 learning rate, batch size 1024, and 300 epochs for DeiT-Tiny/Small. For fine-grained datasets (CUB200, FGVC-Aircrafts, Stanford-Cars), we fine-tune DeiT-Small and ViT-Base for 200 epochs, batch size 16, learning rate 1e-5, using PyTorch pretrained weights (Paszke et al., 2019).
Following LLaVA-v1.5, we adopt Vicuna-v1.5 7B (Chiang et al., 2023) as the language decoder and use a pre-trained 2-layer MLP projection to align visual and textual modalities, trained for one epoch on LCS-558K. The vision encoder is a pre-trained CLIP model that extracts image representations. During SFT, we train for one epoch on llava-v1.5-mix665k with a 2e-5 learning rate, batch size 64, and AdamW optimizer, using a 0.03 warmup ratio and cosine scheduler. Unlike LLaVA, we unfreeze the vision encoder during training. For Qwen2.5-VL-Instruction, we post training within the official checkpoints and the llava-v1.5-mix665k dataset for 0.1 epoch under similar optimization settings. The learning rates for the vision encoder, LLM decoder, and merger are set to 2e-6, 1e-5, and 1e-5, respectively.
@@ 5.4 Results of Image Classification
We did the experiments on three classification datasets on a small-scale dataset (CIFAR100), a large-scale dataset (ImageNet-1K), and a fine-grained dataset (Stanford-Cars). (1) CIFAR100: Table 1 and Table 9 shows the part and full classification results respectively. MergeMix brings the +2.15%, +2.51% gains compared with TranMix on DeiT models. Gains +2.87%, +2.88% and +4.79% on ViT models. All the results in Table 1 are trained for 200 epochs on the CIFAR100 dataset. (2) Stanford-Cars: Table 2 shows the fine-grain classification results on Stanford-Cars dataset. MergeMix achieves the 88.42% and 92.20% accuracy compared with other mixup methods. About the results of the CUB200 dataset and the FGVC-Aircrafts dataset in Table 10. (3) ImageNet-1k: Table 3 shows the results of accuracy, throughput, and flops on the ImageNet-1K dataset. It is notable that MergeMix brings a +0.27% gains and reduces -0.68G Flops compared with TransMix, and can also see other mixups with less throughput since they bring extra cost, but MergeMix has a high throughput of 1591.66 TP/s.
@@
Figure 5: Robustness against image occlusion classification results with different occlusion ratios for different mixup methods based on DeiT-Small on ImageNet-1K dataset.
@@ 5.5 Results of MLLM Benchmarks
We chose two mainstream MLLMs for our experiments on the VQA tasks and reasoning. Table 4 shows the results of LLaVA benchmarks, with different vision tokens for training. With the setting of full vision tokens on the training stage, our method achieves an average gain of +0.83%. When reducing the vision token to 288, our method still retains good performance compared with SFT. A full comparison of results in Table 16, Table 17, and Table 18. In those different settings, MergeMix can achieve the average performance of 66.84%, which gains +1.27% over the LLaVA baseline. Table 5 shows the results of some VQA tasks and reasoning with Qwen2.5-VL-Instruction. MergeMix achieves an average gain of +2.88% over Qwen2.5-VL-Instruction. For the MathVista reasoning task, we reported the results without LLM-as-the-judge-eval. For the MMMU and MMMU-Pro tasks, we can achieve results on par with methods targeting reasoning improvements.
@@ 5.6 Results of Calibration
DNNs are prone to overconfidence in classification tasks. Mainfold Mixup (Verma et al., 2019) found that the mixup methods can effectively alleviate this problem. To this end, we compute the Expected Calibration Error (ECE (Guo et al., 2017)) of various mixup approaches on the CIFAR100 dataset for image classification. Also, to further analyze the calibration of MLLMs, we implement four short answer tasks, POPE, GQA, ScienceVQA, and SEED. Figure 4 shows the results of DeiT-Tiny and ViT-Small models trained for 200 epochs, showing that MergeMix obtains the best calibration of 6.7% and 9.7% in those ViT-specific mixup methods. (i.e., TransMix, SMMix, and MixPro). Table 6 shows the results of the LLaVA baseline, LLaVA with SFT, and LLaVA with our approach. SFT reduces the ECE when tuning the vision encoder, with augmentation and ranking loss, which can be better since we bring in the reward signal for the model. The more comprehensive results of CIFAR100 and MLLM benchmark we plot in Table 11 and Table 15.
5.7 Results of Occlusion Robustness
To analyze the improvement of MergeMix on the robustness with other mixup methods, we randomly conducted a masking patches experiment (Naseer et al., 2021). The experiment involved dividing the image into 16 $\times$ 16 patches and masking them according to different ratios. The resulting image was then classified by the model. Figure 5 shows the accuracy curve of the ImageNet-1K dataset evaluation by a trained DeiT-Small model. About the full results, we put them in Appendix: Figure 7, Table 12, Table 14, and Table 13.
@@ Appendix A Declaration of LLM Usage
Appendix B Appendix for Image Classification
B.1 Information on Datasets
@@
We choose five mainstream classification datasets: (i) CIFAR100 dataset (Krizhevsky et al., 2009) consists of 100 classes of color images with a resolution of 32 $\times$ 32 pixels, containing 50,000 training images and 10,000 test images. (ii) ImageNet-1k dataset (Russakovsky et al., 2015) consists of 1,000 classes with varied image resolutions commonly cropped to 224 $\times$ 224 pixels, containing 1,281,167 training images and 50,000 test images. (iii) CUB200 dataset (Wah et al., 2011) contains 200 bird species, including total 11,788 images, we divided 5,994 images as training images and 5,794 images as test images, (iv) FGVC-Aircrafts dataset (Maji et al., 2013) contains 100 aircraft model classes, including 6,667 training images and 3,333 test images, (v) Stanford-Cars dataset (Krause et al., 2013) contains 196 car model classes, including 8,144 training images and 8,041 test images. All the fine-grained datasets we used set the resolution as 224 $\times$ 224 pixels for training and testing.
@@ B.2 Detailed Implementations
(i) For the CIFAR100 dataset, aiming to be suitable for training ViT-based approaches, e.g., DeiT, we resize images to $224\times 224$ and train them with the AdamW optimizer with weight decay of 0.05, batch size of 100, and total training of 200 epochs and 600 epochs. Uses RandomFlip and RandomCrop as basic augmentations, and additionally, we use RandAugment (Cubuk et al., 2020). For DeiT-Tiny and DeiT-Small, we use the learning rate of 1e-3 with a dynamic cosine scheduler. For ViT-Small and ViT-Base models, we set the learning rate to 5e-4, the learning rate of ViT-Large up to 2e-4, all dynamically adjusted by a cosine scheduler. (ii) For the ImageNet-1K dataset, the dataset settings are the same as CIFAR100, but we use the learning rate as 1e-3, batch size of 1024, and a total training of 300 epochs for DeiT-Tiny and DeiT-Small with AdamW optimizer with weight decay of 0.05. (iii) For all fine-grain datasets, i.e., CUB-200 dataset, FGVC-Aircrafts dataset, and Stanford-Cars dataset, we fine-tune the DeiT-Small and ViT-Base model for 200 epochs with a batch size of 16, learning rate of 1e-5, loading the pre-trained model weight from PyTorch (Paszke et al., 2019).
Table 9: Top-1 accuracy (%) of mixup methods on CIFAR-100 dataset under DeiT-Tiny/Small, ViT-Small/Base/Large different model sizes. The $\alpha$ parameter of the Beta distribution follows the setting in OpenMixup (Li et al., 2022) setting.
@@
Table 10: Top-1 accuracy (%) of mixup methods on Fine-Grained datasets: CUB200, FGVC-Aircrafts, and Stanford-Cars.
@@
Following the LLaVA-v1.5 settings, we use a pre-trained Vicuna-v1.5 7B (Chiang et al., 2023) as the language decoder, which uses a pre-trained 2 $\times$ MLP as the projection for aligning the vision and text modistes, which was trained for one epoch on LCS-558K. For the vision encoder, we use a pre-trained CLIP encoder and extract the visual representation from the input images. For SFT, the learning rate was set as 2e-5, the batch size was 64, and training one epoch on llava-v1.5-mix665k dataset, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8, warmup ratio of 0.03 with a cosine scheduler. The difference from LLaVA is that we unfreeze the vision encoder during training. About Qwen2.5-VL-Instruction, we fine-tune with llava-v1.5-mix665k dataset for 0.1 epoch, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8 like LLaVA, warmup with 0.03 ratio. The learning rate of the vision encoder, LLM decoder, and merger were set to 2e-6, 1e-5, and 1e-5, respectively.
B.3 Additional Results of Image Classification

### v2

@@ 5 Experiments
To evaluate the performance of MergeMix, we compared with some mainstream mixup methods, i.e. Mixup (zhang2017mixup), CutMix (yun2019cutmix), FMix (harris2020fmix), SmoothMix (lee2020smoothmix), GridMix (Baek2021gridmix), ResizeMix (qin2020resizemix), SaliencyMix (uddinsaliencymix), Attentive-CutMix (icassp2020attentive), PuzzleMix (kim2020puzzle), GuidedMixup (kang2023guidedmixup), AutoMix (liu2022automix) and AdAutoMix (qin2024adversarial). DeiT (icml2021deit), TransMix (chen2022transmix), SMMix (chen2023smmix), MixPro (zhao2023mixpro) and TdAttenMix (wang2025tdattenmix) for some ViT-based methods. The training configures about datasets and methods follows the open-source library OpenMixup (2022openmixup).
@@
To evaluate the training paradigm that we proposed, we compare with three different system-level baselines: (1) SFT with different training paradigms on LLaVA, including LLaVA-NeXT-7/13B (liu2024llavanext), SeVa-7B (zhu2024seva), SIMA (wang2024sima), and nSFT (zhu2025nsft). (2). Token reduction on LLaVA, including LLaVA-PruMerge+ (shang2024llavapruerge), VisionZip (yang2025visionzip), VisPrunner (zhang2024vispunner), VScan (zhang2025vscan), and LLaVA-Mini (zhang2025llava-mini). (3). RL training on Qwen2.5-VL-Instruction (bai2025qwen2.5), including VisionThink (yang2025visionthink).
For all classification results, we report the top-1 test accuracy in the last 10 training epochs for each trial. To facilitate comparison, we mark the best and second best results in bold and cyan. For the LLaVA benchmark, we use the LLaVA official code, and for the Qwen2.5-VL-Instruction benchmark. We use lmms-eval (zhang2024lmmsevalrealitycheckevaluation) for evaluation.
Table 1: Top-1 accuracy (%) of mixup methods training 200 epochs on CIFAR100 dataset with different model sizes, T/S/B/L denotes Tiny, Small, Base, and Large, respectively. The full results of training 600 epochs are in Table A1.
@@
Table 2: Top-1 accuracy (%) of mixup methods on the Stanford-Cars dataset. Full results of the CUB200 and FGVC-Aircrafts dataset in Table A2.
@@ 5.2 Datasets
In our paper, we mainly divided into 2 scenarios: Image Classification and MLLM Benchmark. The detailed information about datasets is described in Appendix B.1. For the image classification datasets, we choose 5 public classification datasets, including the small-scale dataset of CIFAR100 (krizhevsky2009learning), the large-scale dataset of ImageNet-1K (russakovsky2015imagenet), and the fine-grained datasets of CUB200 dataset (wah2011caltech), FGVC-Aircrafts dataset (maji2013fine), and Stanford-Cars dataset (jonathan2013cars). For the MLLM datasets, we choose 16 datasets, including visual question answering (VQAv2 (goyal2017vqav2), GQA (hudson2019gqa), VizWiz (gurari2018vizwiz), ScienceVQAI (lu2022scivqa), TextVQA (singh2019textvqa), MME-RealWorldQA (zhang2025realworldqa)), understanding (MME (Perception) (yin2023mme), MMBench (liu2025mmbench), $^{\text{CN}}, MMBench$$^{\text{CC}}$, POPE (F1 score) (li2023pope), SEEDI (li2023seed), MMStar (chen2024mmstar)), and reasoning (MMMU (yue2024mmmu), MMMU-Pro Standard (MMMU-Pros) (yue2024mmmu-pro), MathVista (lu2024mathvista)).
Table 4: Full system-level comparison results in LLaVA. Compared with their counterparts. AVG denotes the average of the nine benchmarks for comprehensive comparison, except for MME, underline denotes MME with the sum of Perception and Cognition. Tokeni denotes training with the token number. Full results in Table A10, Table A11 and Table A12.
@@ 5.3 Implementations
In this subsection, we briefly introduce the implementations on the classification task and the MLLM benchmark. The full description is in Appendix B.2. For the CIFAR100, images are resized to $224\times 224$ for ViT-based models (e.g., DeiT) and trained with AdamW (weight decay 0.05), batch size 100, for 200 or 600 epochs. We use RandomFlip and RandomCrop, plus RandAugment (cubuk2020randaugment). Learning rates are 1e-3 (DeiT-Tiny/Small, cosine schedule), 5e-4 (ViT-Small/Base), and 2e-4 (ViT-Large). For ImageNet-1K, we adopt the same settings but use a 1e-3 learning rate, batch size 1024, and 300 epochs for DeiT-Tiny/Small. For fine-grained datasets (CUB200, FGVC-Aircrafts, Stanford-Cars), we fine-tune DeiT-Small and ViT-Base for 200 epochs, batch size 16, learning rate 1e-5, using PyTorch pretrained weights (nips2019pytorch).
Following LLaVA-v1.5, we adopt Vicuna-v1.5 7B (chiang2023vicuna) as the language decoder and use a pre-trained 2-layer MLP projection to align visual and textual modalities, trained for one epoch on LCS-558K. The vision encoder is a pre-trained CLIP model that extracts image representations. During SFT, we train for one epoch on llava-v1.5-mix665k with a 2e-5 learning rate, batch size 64, and AdamW optimizer, using a 0.03 warmup ratio and cosine scheduler. Unlike LLaVA, we unfreeze the vision encoder during training. For Qwen2.5-VL-Instruction, we post training within the official checkpoints and the llava-v1.5-mix665k dataset for 0.1 epoch under similar optimization settings. The learning rates for the vision encoder, LLM decoder, and merger are set to 2e-6, 1e-5, and 1e-5, respectively.
@@ 5.4 Results of Image Classification
We did the experiments on three classification datasets on a small-scale dataset (CIFAR100), a large-scale dataset (ImageNet-1K), and a fine-grained dataset (Stanford-Cars). (1) CIFAR100: Table 1 and Table A1 shows the part and full classification results respectively. MergeMix brings the +2.15%, +2.51% gains compared with TranMix on DeiT models. Gains +2.87%, +2.88% and +4.79% on ViT models. All the results in Table 1 are trained for 200 epochs on the CIFAR100 dataset. (2) Stanford-Cars: Table 2 shows the fine-grain classification results on Stanford-Cars dataset. MergeMix achieves the 88.42% and 92.20% accuracy compared with other mixup methods. About the results of the CUB200 dataset and the FGVC-Aircrafts dataset in Table A2. (3) ImageNet-1k: Table 3 shows the results of accuracy, throughput, and flops on the ImageNet-1K dataset. It is notable that MergeMix brings a +0.27% gains and reduces -0.68G Flops compared with TransMix, and can also see other mixups with less throughput since they bring extra cost, but MergeMix has a high throughput of 1591.66 TP/s.
@@
Figure 5: Visualization of MergeMix on different mixing ratios, including mixed images, GradCAM of top-2 logits, and ToMe source maps.
@@ 5.5 Results of MLLM Benchmarks
We chose two mainstream MLLMs for our experiments on the VQA tasks and reasoning. Table 4 shows the results of LLaVA benchmarks, with different vision tokens for training. With the setting of full vision tokens on the training stage, our method achieves an average gain of +0.83%. When reducing the vision token to 288, our method still performs well compared with SFT. A full comparison of results in Table A10, Table A11, and Table A12. In those different settings, MergeMix achieves an average performance of 66.84%, improving over the LLaVA baseline by +1.27%. Table 5 shows the results of some VQA tasks and reasoning with Qwen2.5-VL-Instruction. MergeMix achieves an average gain of +2.88% over Qwen2.5-VL-Instruction. For the MathVista reasoning task, we reported the results without LLM-as-the-judge-eval. For the MMMU and MMMU-Pro tasks, we can achieve results on par with methods targeting reasoning improvements.
@@ 5.6 Results of Calibration
DNNs are prone to overconfidence in classification tasks. Mainfold Mixup (verma2019manifold) found that the mixup methods can effectively alleviate this problem. To this end, we compute the Expected Calibration Error (ECE (guo2017calibration)) of various mixup approaches on the CIFAR100 dataset for image classification. Also, to further analyze the calibration of MLLMs, we implement four short answer tasks, POPE, GQA, ScienceVQA, and SEED. Figure 4 shows the results of DeiT-Tiny and ViT-Small models trained for 200 epochs, showing that MergeMix obtains the best calibration of 6.7% and 9.7% in those ViT-specific mixup methods. (i.e., TransMix, SMMix, and MixPro). Table 6 shows the results of the LLaVA baseline, LLaVA with SFT, and LLaVA with our approach. SFT reduces the ECE when tuning the vision encoder, with augmentation and ranking loss, which can be better since we bring in the reward signal for the model. The more comprehensive results of CIFAR100 and MLLM benchmark we plot in Table A3 and Table A9.
@@ 6 Conclusion
There remain limitations in MergeMix for MLLMs. In future work, we will address them from two directions: (1) Data level: MergeMix currently enhances only the image modality, while text remains untouched. Extending mixup to the text modality could provide more fine-grained optimization. (2) Model level: The token-merging policy is static and unlearned. Making it learnable may further improve the mixing capability.
@@ Appendix A Declaration of LLM Usage
Appendix B Detail Information
B.1 Datasets
@@
We choose five mainstream classification datasets: (i) CIFAR100 dataset (krizhevsky2009learning) consists of 100 classes of color images with a resolution of 32 $\times$ 32 pixels, containing 50,000 training images and 10,000 test images. (ii) ImageNet-1k dataset (russakovsky2015imagenet) consists of 1,000 classes with varied image resolutions commonly cropped to 224 $\times$ 224 pixels, containing 1,281,167 training images and 50,000 test images. (iii) CUB200 dataset (wah2011caltech) contains 200 bird species, including total 11,788 images, we divided 5,994 images as training images and 5,794 images as test images, (iv) FGVC-Aircrafts dataset (maji2013fine) contains 100 aircraft model classes, including 6,667 training images and 3,333 test images, (v) Stanford-Cars dataset (jonathan2013cars) contains 196 car model classes, including 8,144 training images and 8,041 test images. All the fine-grained datasets we used set the resolution as 224 $\times$ 224 pixels for training and testing.
@@ B.2 Implementations
(i) For the CIFAR100 dataset, aiming to be suitable for training ViT-based approaches, e.g., DeiT, we resize images to $224\times 224$ and train them with the AdamW optimizer with weight decay of 0.05, batch size of 100, and total training of 200 epochs and 600 epochs. Uses RandomFlip and RandomCrop as basic augmentations, and additionally, we use RandAugment (cubuk2020randaugment). For DeiT-Tiny and DeiT-Small, we use the learning rate of 1e-3 with a dynamic cosine scheduler. For ViT-Small and ViT-Base models, we set the learning rate to 5e-4, the learning rate of ViT-Large up to 2e-4, all dynamically adjusted by a cosine scheduler. (ii) For the ImageNet-1K dataset, the dataset settings are the same as CIFAR100, but we use the learning rate as 1e-3, batch size of 1024, and a total training of 300 epochs for DeiT-Tiny and DeiT-Small with AdamW optimizer with weight decay of 0.05. (iii) For all fine-grain datasets, i.e., CUB-200 dataset, FGVC-Aircrafts dataset, and Stanford-Cars dataset, we fine-tune the DeiT-Small and ViT-Base model for 200 epochs with a batch size of 16, learning rate of 1e-5, loading the pre-trained model weight from PyTorch (nips2019pytorch).
Table A1: Top-1 accuracy (%) of mixup methods on CIFAR-100 dataset under DeiT-Tiny/Small, ViT-Small/Base/Large different model sizes. The $\alpha$ parameter of the Beta distribution follows the setting in OpenMixup (2022openmixup) setting.
@@
Table A2: Top-1 accuracy (%) of mixup methods on Fine-Grained datasets: CUB200, FGVC-Aircrafts, and Stanford-Cars.
@@
Following the LLaVA-v1.5 settings, we use a pre-trained Vicuna-v1.5 7B (chiang2023vicuna) as the language decoder, which uses a pre-trained 2 $\times$ MLP as the projection for aligning the vision and text modistes, which was trained for one epoch on LCS-558K. For the vision encoder, we use a pre-trained CLIP encoder and extract the visual representation from the input images. For SFT, the learning rate was set as 2e-5, the batch size was 64, and training one epoch on llava-v1.5-mix665k dataset, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8, warmup ratio of 0.03 with a cosine scheduler. The difference from LLaVA is that we unfreeze the vision encoder during training. About Qwen2.5-VL-Instruction, we fine-tune with llava-v1.5-mix665k dataset for 0.1 epoch, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8 like LLaVA, warmup with 0.03 ratio. The learning rate of the vision encoder, LLM decoder, and merger were set to 2e-6, 1e-5, and 1e-5, respectively.
B.3 Algorithms
Algorithm 1 and Algorithm 2 show the pseudo codes for MergeMix on both image classification and preference tuning.
Algorithm 1 MergeMix for Image Classification
# Inputs: vision model $f_{\theta}(\cdot)$, training parameters $\theta$ of model, dataset $D$, mixup parameter $\alpha$; output: updated $f_{\theta}(\cdot)$
#  sample a paired mini-batch from $D$ (two images and labels)
1: for ($x_{i},y_{i}$), ($x_{j},y_{j}$) in DataLoader($D$):
#  sample mixup ratio $\lambda$ (shared with Alg. 2)
2:  $\lambda\sim\mathrm{Beta}(\alpha,\alpha)$
3:  $\mathcal{M},\hat{\lambda}=\mathcal{P}(x_{i},x_{j},\lambda)$
#  lines 3--4: MergeMix augmentation (same core as Alg. 2)
4:  $\hat{x}=\mathcal{M}\odot x\_i+(1-\mathcal{M})\odot x\_j$
[…]
#  main supervised loss on mixed sample (parallels $L_{\text{SFT}}$ in Alg. 2)
6:  $\mathcal{L}_{\text{MCE}}=\hat{\lambda}\cdot\mathcal{L}_{\text{CE}}(\textrm{logits},y_{i})+(1-\hat{\lambda})\cdot\mathcal{L}_{\text{CE}}(\textrm{logits},y_{j})$
#  optional regularizer for augmentation consistency
7:  $\mathcal{L}_{\text{CE}}=f_{\theta}(x,y)$
8:  $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{MCE}}+\mathcal{L}_{\text{CE}}$
9:  $\mathcal{L}_{\text{Total}}$.backward()
10: optimizer.step()
11: optimizer.zero_grad()
Algorithm 2 MergeMix for MLLM Alignment
# Inputs: MLLM $\pi_{\theta}(\cdot)$, learnable parameters $\theta$ of model, dataset $D$ of $(x,q,T)$; output: updated $\pi_{\theta}(\cdot)$
#  iterate over triples
1: for $(x,q,T)$ in DataLoader($D$):
#  sample an auxiliary image $x_{i}$ (mirrors mix partner in Alg. 1)
2:  ($x_{i},$\_) = randomSample($D$)
3:  $\mathcal{M}=\mathcal{P}(x,x_{i})$
#  lines 3--4: same MergeMix augmentation core as Alg. 1
4:  $\hat{x}=\mathcal{M}\odot x+(1-\mathcal{M})\odot x_{i}$
#  winner (raw) vs. loser (mixed) outputs for the same prompt
5:  $Y_{w}=\pi_{\theta}(x,q)$
6:  $Y_{l}=\pi_{\theta}(\hat{x},q)$
#  supervised fine-tuning loss (analogous to $\mathcal{L}_{\text{CE}}$ in Alg. 1)
7:  $\mathcal{L}_{\text{SFT}}=\mathcal{L}_{\text{CE}}(Y_{\text{w}},T)$
8:  $s_{w}=\mathrm{AvgLogProb}(Y_{w},T)$
9:  $s_{l}=\mathrm{AvgLogProb}(Y_{l},T)$
#  preference loss (SimPO) that prefers winner over loser
10: $\mathcal{L}^{\text{Mix}}_{\text{SimPO}}=-\log\sigma(s_{w}-s_{l}-\lambda)$
11: $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{SFT}}+\mathcal{L}^{\text{Mix}}_{\text{SimPO}}$
12: $\mathcal{L}_{\text{Total}}$.backward()
13: optimizer.step()
14: optimizer.zero_grad()
B.4 Future work
Appendix C Image Classification
C.1 Extensive Results
Figure A1: Robustness against image occlusion classification results with different occlusion ratios for different mixup methods based on DeiT-Tiny (left) and DeiT-Small (right) on CIFAR100 and ImageNet-1K datasets.